# Search for high energy 5.5 MeV solar axions with the complete Borexino dataset

### v1

@@ Abstract
A search for solar axions and axion-like particles produced in the $p+d\rightarrow\rm{{}^{3}He}+A~{}(5.5\rm{~{}MeV})$ reaction was performed using the complete dataset of the Borexino detector ($3995$ days of measurement live-time). The following interaction processes have been considered: axion decay into two photons $({\rm A}\rightarrow 2\gamma)$, inverse Primakoff conversion on nuclei $({\rm A}+Z\rightarrow\gamma+Z$), the Compton conversion of axions to photons $({\rm A}+e\rightarrow e+\gamma)$ and the axio-electric effect $({\rm A}+e+Z\rightarrow e+Z$). Model-independent limits on axion-photon ($g_{A\gamma}$), axion-electron ($g_{Ae}$), and isovector axion-nucleon ($g_{3AN}$) couplings are obtained: $|g_{A\gamma}\times g_{3AN}|\leq 2.3\times 10^{-11}\rm{GeV}^{-1}$ and $|g_{Ae}\times g_{3AN}|\leq 1.9\times 10^{-13}$ at $m_{A}<$ 1 MeV (90% c.l.). The Borexino results exclude new large regions of $g_{A\gamma}$, and $g_{Ae}$ coupling constants and axion masses $m_{A}$, and leads to constraints on the products $|g_{A\gamma}\times m_{A}|$ and $|g_{Ae}\times m_{A}|$ for the KSVZ- and the DFSZ-axion models.
Keywords:
axion, pseudoscalar particles; low background measurements
@@ 1 Introduction
The goal of this study is to search for solar axions with an energy of $5.5$ MeV, produced in the $p+d\rightarrow\rm{{}^{3}He}+A$ ($5.49$ MeV) reaction using the large neutrino scintillation detector. The axion flux is thus proportional to the $pp$-neutrino flux, which is known with a high accuracy Ser09 ; Vin17 ; Bel11A ; Ago19 .
@@
We have previously published the result of searches for solar axions emitted in the $478$ keV M1-transition of ${}^{7}{\rm{Li}}$ using the Borexino counting test facility Bel08 and $5.5$ MeV solar axions using 536$\times$100 days×$\times$×tons of Borexino detector data Bel12a .
@@
The value of $5.5$ MeV solar axion flux can be expressed in terms of the $pp$-neutrino flux because $99.7$% of deuterium is produced from the fusion of two protons reaction: $p+p\rightarrow d+e^{+}+\nu_{e}$ The proportionality factor between the axion and $pp$-neutrino fluxes is determined by a ratio between the probability of nuclear transition with axion production $(\omega_{A})$ and photon production $(\omega_{\gamma})$. The ratio $(\omega_{A}/\omega\gamma)$ depends on dimensionless axion-nucleon coupling constant $g_{AN}$, which consists of isoscalar $g_{0AN}$ and isovector $g_{3AN}$ components. The isoscalar and isovector parts of the axion–nucleon coupling constant, respectively, are model dependent. They can be expressed in terms of the effective axion–proton and axion–neutron coupling constants $C_{p}$ and $C_{n}$, respectively Cor16 ; Avi18 :
@@
In the DFSZ-model $C_{p}$ and $C_{n}$ depend on an additional parameter $\beta$ which is the ratio of the vacuum expectation values of the two Higgs doublets giving masses to the up- and down-quarks: $\rm{tan}\beta=\upsilon_{u}/\upsilon_{d}$. The values $C^{DFSZ}_{p}=-(0.435~{}\rm{sin^{2}}\beta+0.182)\pm 0.025$ and $C^{DFSZ}_{n}=(0.414~{}\rm{sin^{2}}\beta-0.160)\pm 0.025$ for the DFSZ axion were obtained in Cor16 in terms of the corresponding model-dependent quark couplings $C_{u}=C_{c}=C_{t}=1/3~{}\rm{cos}^{2}\beta$ and $C_{d}=C_{s}=C_{b}=1/3~{}\rm{sin}^{2}\beta$. The possible values of $\rm{tan}\beta$ are restricted by the range $0.25\leq\rm{tan}\beta\leq 140$ Che13 ; Luz20 . Using Eqs. (1, 3, 4), one can express isoscalar $g_{0AN}$ and isovector $g_{3AN}$ constants in terms of the axion mass $m_{A}$.
Figure 1: The expected fluxes of the solar axions emitted in the $p+d\rightarrow{{}^{3}{\rm{He}}}+A$ reaction. 1) KSVZ, 2) DFSZ ($\rm{tan}\beta=140$), 3)DFSZ ($\rm{tan}\beta=0.25$), 4, 5, 6) - the decay $A\rightarrow 2\gamma$ is taken into account, in the KSVZ- and DFSZ-models, respectively. At $m_{A}\geq 2m_{e}$ the decay $A\rightarrow e^{+}e^{-}$ is possible (dot line).
The M1-type transition in the $p+d\rightarrow\rm{{}^{3}He}+\gamma$ reaction corresponds to the capture of a proton from the $S$ state. The probability, $\chi$, of proton capture with zero orbital moment at energies below 80 keV was measured in Sch97 ; at a proton energy of $\sim 1$ keV, $\chi$ = 0.55. The proton capture from the $S$ state corresponds is an isovector transition, and the ratio $\omega_{A}/\omega_{\gamma}$ depends only on $g_{3AN}$ Don78 ; Hax91 ; Avi88 :
$\frac{\omega_{A}}{\omega_{\gamma}}=\frac{\chi}{2\pi\alpha}\left[\frac{g_{3AN}}{\mu_{3}}\right]^{2}\left(\frac{p_{A}}{p_{\gamma}}\right)^{3}=0.54(g_{3AN})^{2}\left(\frac{p_{A}}{p_{\gamma}}\right)^{3}.$ (5)
where $p_{\gamma}$ and $p_{A}$ are, respectively, the photon and axion momenta, $\alpha\approx 1/137$ is the fine-structure constant, $\mu_{3}=\mu_{p}-\mu_{n}\approx 4.71\mu_{N}$ is the isovector nuclear magnetic moment in units of nuclear magneton and $g^{KSVZ}_{3AN}=-0.49~{}(M_{N}/2f_{A})$ and $g^{DFSZ}_{3AN}=-(M_{N}/2f_{A})(0.849~{}\rm{sin^{2}}\beta+0.022)$ for KSVZ- and DSVZ-models, respectively.
@@
where $\Phi_{\nu pp}=6.0\times 10^{10}{\rm{cm}}^{-2}{\rm{s}}^{-1}$ is the $pp$ solar neutrino flux Ser09 ; Vin17 ; Bel11A ; Ago19 . Using the relation between $g_{3AN}$ and $m_{A}$ given by (4), the $\Phi_{A0}$ value appears to be proportional to $m_{A}^{2}$: $\Phi_{A0}=5.28\times 10^{-5}m_{A}^{2}(p_{A}/p_{\gamma})^{3}~{}\rm{cm^{-2}~{}s^{-1}}$ (KSVZ), $\Phi_{A0}=16.7\times 10^{-5}m_{A}^{2}(p_{A}/p_{\gamma})^{3}~{}\rm{cm^{-2}~{}s^{-1}}$ (DFSZ, $\rm{tan}\beta=140$), and $\Phi_{A0}=1.14\times 10^{-6}m_{A}^{2}(p_{A}/p_{\gamma})^{3}~{}\rm{cm^{-2}~{}s^{-1}}$ (DFSZ, $\rm{tan}\beta=0.25$), where $m_{A}$ is given in eV units. Apparently, the DFSZ-axion flux can be more than two orders of magnitude lower for certain values of the $\beta$-parameter (see Fig. 1).
The dependence of the flux $\Phi_{A0}$ on the axion mass $m_{A}$ has a bell-shaped form reaching a maximum at an energy of $3.5$ MeV, for small $m_{A}$ the flux is suppressed by the factor $m_{A}^{2}$ and for large $m_{A}$ by the factor $(p_{A}/p_{\gamma})^{3}$. The calculated values of the uncharged fluxes $\Phi_{A0}(m_{A})$ as a function of the axion mass are shown in Fig. 1 for KSVZ- and DSVZ-models (lines 1, 2 and 3).
@@
where $E/N$ is a model dependent parameter of the order of unity. $E/N=8/3$ in the DFSZ axion models ($C_{A\gamma\gamma}$=0.74) and $E/N=0$ for the original KSVZ axion ($C_{A\gamma\gamma}=-1.92$). Really, there are QCD axion models whose $g_{A\gamma}$ couplings constants are outside the region restricted by this two values, motivating searches for axions over a wide range of masses and couplings constants.
The decay $\rm{A}\rightarrow 2\gamma$ was the first reaction used for experimental search for standard PQWW-axion. The probability of the decay is given by the expression:
$\tau_{2\gamma}=\frac{64\pi}{g_{A\gamma}^{2}m_{A}^{3}}.$ (8)
@@
$\tau_{2\gamma}=1.3\times 10^{5}g_{A\gamma}^{-2}m_{A}^{-3}=3.5\times 10^{24}m_{A}^{-5}C_{A\gamma\gamma}^{-2}.$ (9)
@@
where $\Phi_{A0}$ is the axion flux at the Earth in case there is no axion decay (6), $\tau_{2\gamma}$ is defined by (8, 9), and $\tau_{f}$ is the time of flight in the axion frame of reference. Because of axion decay, the sensitivity of experiments using solar axions drops off for large values of $g_{A\gamma}^{2}m_{A}^{3}$. The expected axion spectra for KSVZ- and DFSZ-models as function of $m_{A}$ are shown in Fig. 1 (lines 4 KSVZ-, and lines 5 and 6 for DFSZ-axions). The fluxes have a bell-shaped form with maximum at $\sim(40-50)$ keV since at low $m_{A}$ the probability of decay is low and at large ones the axion flux is reduced due to decays $\rm{A}\rightarrow 2\gamma$ on the way to the Earth.
The results were interpreted in terms of QCD-axion models using relevant relations between the axion mass and the coupling constants. In case of axion-like particles the QCD-axion relations no longer apply and the results must be analyzed with axion mass and couplings as independent parameters.
@@ 3.2 Compton conversion $A+e\rightarrow\gamma+e$ and axio-electric effect $A+e+Z\rightarrow e+Z$
Two others studied reactions - the Compton process ${\rm A}+e\rightarrow e+\gamma$, and the axioelectric effect ${\rm A}+e+Z\rightarrow e+Z$ depend on the coupling constant between the axion and the electron $g_{Ae}$.
The dimensionless coupling constant $g_{Ae}$ is associated with the electron mass $m_{e}$, so that $g_{Ae}=C_{e}m_{e}/f_{A}$, where $C_{e}$ is a model dependent factor of the order of unity. In the standard PQWW axion model, the values $f_{A}$=250 GeV and $C_{e}$=1 are fixed and $g_{Ae}\approx 2\times 10^{-6}$. In the DFSZ axion models $C_{e}=\sin^{2}\beta/3$, where as was mentioned above $\beta$ is an unknown angle, such that $\tan\beta=\upsilon_{u}/\upsilon_{d}$ is the ratio of the vacuum values of the two Higgs doublets. Assuming e.g. $\tan\beta$=10 Avi18 , the axion-electron coupling is $g_{Ae}$=2.96×$\times 10^{-11}m_{A}$ where $m_{A}$ is expressed in $eV$ units.
@@
$g_{Ae}=\frac{3\alpha^{2}m_{e}}{4\pi^{2}f_{A}}\left(\frac{E}{N}\ln\frac{f_{A}}{m_{e}}-\frac{2}{3}\frac{4+z}{1+z}\ln\frac{\Lambda}{m_{e}}\right)$ (12)
@@
The axio-electric effect $A+e+Z\rightarrow e+Z$ is the analogue of the photo-electric effect and depends on $g_{Ae}$-coupling. The cross-sections of the axioelectric effect, expressed through the cross-section of the photo-effect, including for the case of non-relativistic axions, were obtained in work Pos08 . The cross section of the axio-electric effect on K-electrons where the axion energy $E_{A}\gg E_{b}$ was calculated in Zhi79 and has a complex form also. The cross section has a $Z^{5}$ dependence and for carbon atoms the cross section is $\sigma_{Ae}\approx g_{Ae}^{2}\times$1.3$\times$10-29 cm2/electron for $m_{A}<$ 1 MeV.
@@ 4.1 Borexino detector
The design and main components and features of the Borexino is described in detail in Ago19 ; Ali02 ; Ali09 ; Bac12 ; Bel14 . Borexino is a scintillator detector with an active mass of 278 tons of pseudocumene, doped with $1.5$ g/L of PPO. The scintillator is housed in a thin nylon inner vessel (IV) and is surrounded by two concentric pseudocumene buffers ($323$ and $567$ tons) doped with a small amount of light quencher to reduce their scintillation. The two buffers are separated by a second thin nylon membrane to prevent diffusion of external radon. The scintillator and buffers are contained in a stainless steel sphere (SSS) with diameter $13.7$ m. The SSS is enclosed in domed water tank (WT), containing 2100 tons of ultra pure water as a shield against external $\gamma$’s and neutrons. The scintillation light is detected by $2212$ $8^{\prime\prime}$ PMTs distributed on the SSS. The WT is equipped with $208$ additional PMTs that act as a Cerenkov muon detector (outer detector) to identify cosmogenic muons. The energy and spatial resolution of the detector were studied with radioactive sources placed at different points inside the IV Bel10 ; Bel10A ; Bel11d .

### v2

@@ Abstract
A search for solar axions and axion-like particles produced in the $p+d\rightarrow\rm{{}^{3}He}+A~(5.5\rm{~MeV})$ reaction was performed using the complete dataset of the Borexino detector ($3995$ days of measurement live-time). The following interaction processes have been considered: axion decay into two photons $({\rm A}\rightarrow 2\gamma)$, inverse Primakoff conversion on nuclei $({\rm A}+Z\rightarrow\gamma+Z$), the Compton conversion of axions to photons $({\rm A}+e\rightarrow e+\gamma)$ and the axio-electric effect $({\rm A}+e+Z\rightarrow e+Z$). Model-independent limits on product of axion-photon ($g_{A\gamma}$), axion-electron ($g_{Ae}$), and isovector axion-nucleon ($g_{3AN}$) couplings are obtained: $|g_{A\gamma}\times g_{3AN}|\leq 2.3\times 10^{-11}\rm{GeV}^{-1}$ and $|g_{Ae}\times g_{3AN}|\leq 1.9\times 10^{-13}$ at $m_{A}<$ 1 MeV (90% c.l.). The Borexino results exclude new large regions of $g_{A\gamma}$, and $g_{Ae}$ coupling constants and axion masses $m_{A}$, and leads to constraints on the products $|g_{A\gamma}\times m_{A}|$ and $|g_{Ae}\times m_{A}|$ for the KSVZ- and the DFSZ-axion models.
@@ 1 Introduction
The goal of this study is to search for solar axions with an energy of $5.5$ MeV, produced in the $p+d\rightarrow\rm{{}^{3}He}+A$ ($5.49$ MeV) reaction using the large neutrino scintillation the Borexino detector. The axion flux is thus proportional to the $pp$-neutrino flux, which is known with a high accuracy Ser09 ; Vin17 ; Bel11A ; Ago19 .
@@
[…]
@@
The value of $5.5$ MeV solar axion flux can be expressed in terms of the $pp$-neutrino flux because 99.799.7% of deuterium is produced from the fusion of two protons reaction: $p+p\rightarrow d+e^{+}+\nu_{e}$ . The proportionality factor between the axion and $pp$-neutrino fluxes is determined by a ratio between the probability of nuclear transition with axion production $(\omega_{A})$ and photon production $(\omega_{\gamma})$. The ratio $(\omega_{A}/\omega\gamma)$ depends on dimensionless axion-nucleon coupling constant $g_{AN}$, which consists of isoscalar $g_{0AN}$ and isovector $g_{3AN}$ components. The isoscalar and isovector parts of the axion–nucleon coupling constant, respectively, are model dependent. They can be expressed in terms of the effective axion–proton and axion–neutron coupling constants $C_{p}$ and $C_{n}$, respectively Cor16 ; Avi18 :
@@
In the DFSZ-model, $C_{p}$ and $C_{n}$ depend on an additional parameter $\beta^{*}$ which is the ratio of the vacuum expectation values of the two Higgs doublets giving masses to the up- and down-quarks: $\rm{tan}\beta^{*}=\upsilon_{u}/\upsilon_{d}$. The values $C^{DFSZ}_{p}=-(0.435~\rm{sin^{2}}\beta^{*}+0.182)\pm 0.025$ and $C^{DFSZ}_{n}=(0.414~\rm{sin^{2}}\beta^{*}-0.160)\pm 0.025$ for the DFSZ axion were obtained in Cor16 in terms of the corresponding model-dependent quark couplings $C_{u}=C_{c}=C_{t}=1/3~\rm{cos}^{2}\beta^{*}$ and $C_{d}=C_{s}=C_{b}=1/3~\rm{sin}^{2}\beta^{*}$. The possible values of $\rm{tan}\beta^{*}$ are restricted by the range $0.25\leq\rm{tan}\beta^{*}\leq 140$ Che13 ; Luz20 . Using Eqs. (1, 3, 4), one can express isoscalar $g_{0AN}$ and isovector $g_{3AN}$ constants in terms of the axion mass $m_{A}$.
Figure 1: The expected fluxes of the solar axions emitted in the $p+d\rightarrow{{}^{3}{\rm{He}}}+A$ reaction. 1 - KSVZ, 2 - DFSZ ($\rm{tan}\beta^{*}=140$), 3 - DFSZ ($\rm{tan}\beta^{*}=0.25$), 4, 5, 6 - the decay $A\rightarrow 2\gamma$ is taken into account, in the KSVZ- and DFSZ-models, respectively. At $m_{A}\geq 2m_{e}$ the decay $A\rightarrow e^{+}e^{-}$ is possible (dot line).
The M1-type transition in the $p+d\rightarrow\rm{{}^{3}He}+\gamma$ reaction corresponds to the capture of a proton from the $S$ state. The probability, $\chi$, of proton capture with zero orbital moment at energies below 80 keV was measured in Sch97 ; at a proton energy of $\sim 1$ keV, $\chi$ = 0.55. The proton capture from the $S$ state corresponds an isovector transition, and the ratio $\omega_{A}/\omega_{\gamma}$ depends only on $g_{3AN}$ Don78 ; Hax91 ; Avi88 :
$\frac{\omega_{A}}{\omega_{\gamma}}=\frac{\chi}{2\pi\alpha}\left[\frac{g_{3AN}}{\mu_{3}}\right]^{2}\left(\frac{p_{A}}{p_{\gamma}}\right)^{3}=0.54(g_{3AN})^{2}\left(\frac{p_{A}}{p_{\gamma}}\right)^{3},$ (5)
where $p_{\gamma}$ and $p_{A}$ are, respectively, the photon and axion momenta, $\alpha\approx 1/137$ is the fine-structure constant, $\mu_{3}=\mu_{p}-\mu_{n}\approx 4.71\mu_{N}$ is the isovector nuclear magnetic moment in units of nuclear magneton and $g^{KSVZ}_{3AN}=-0.49~(M_{N}/2f_{A})$ and $g^{DFSZ}_{3AN}=-(M_{N}/2f_{A})(0.849~\rm{sin^{2}}\beta^{*}+0.022)$ for KSVZ- and DSVZ-models, respectively.
@@
where $\Phi_{\nu pp}=6.0\times 10^{10}{\rm{cm}}^{-2}{\rm{s}}^{-1}$ is the $pp$ solar neutrino flux Ser09 ; Vin17 ; Bel11A ; Ago19 . Using the relation between $g_{3AN}$ and $m_{A}$ given by (4), the $\Phi_{A0}$ value appears to be proportional to $m_{A}^{2}$: $\Phi_{A0}=5.28\times 10^{-5}m_{A}^{2}(p_{A}/p_{\gamma})^{3}~\rm{cm^{-2}~s^{-1}}$ (KSVZ), $\Phi_{A0}=16.7\times 10^{-5}m_{A}^{2}(p_{A}/p_{\gamma})^{3}~\rm{cm^{-2}~s^{-1}}$ (DFSZ, $\rm{tan}\beta^{*}=140$), and $\Phi_{A0}=1.14\times 10^{-6}m_{A}^{2}(p_{A}/p_{\gamma})^{3}~\rm{cm^{-2}~s^{-1}}$ (DFSZ, $\rm{tan}\beta^{*}=0.25$), where $m_{A}$ is given in eV units. Apparently, the DFSZ-axion flux can be more than two orders of magnitude lower for certain values of the $\beta^{*}$-parameter (see Fig. 1).
The dependence of the flux $\Phi_{A0}$ on the axion mass $m_{A}$ has a bell-shaped form reaching a maximum at an energy of $3.5$ MeV, for small $m_{A}$ the flux is suppressed by the factor $m_{A}^{2}$ and for large $m_{A}$ by the factor $(p_{A}/p_{\gamma})^{3}$. The calculated values of the uncharged fluxes $\Phi_{A0}(m_{A})$ as a function of the axion mass are shown in Fig. 1 for KSVZ- and DSVZ-models (lines 1, 2, and 3).
@@
where $E/N$ is a model dependent parameter of the order of unity. $E/N=8/3$ in the DFSZ axion models ($C_{A\gamma\gamma}$=0.74) and $E/N=0$ for the original KSVZ axion ($C_{A\gamma\gamma}=-1.92$). There are QCD axion models whose $g_{A\gamma}$ couplings constant are outside the region restricted by this two values, motivating searches for axions over a wide range of masses and couplings constants.
The decay $\rm{A}\rightarrow 2\gamma$ was the first reaction used for experimental search for standard PQWW-axion. The mean lifetime of the decay is given by the expression:
$\tau_{2\gamma}=\frac{64\pi\hbar}{g_{A\gamma}^{2}m_{A}^{3}}.$ (8)
@@
$\tau_{2\gamma}~(s)=1.3\times 10^{5}g_{A\gamma}^{-2}m_{A}^{-3}=3.5\times 10^{24}m_{A}^{-5}C_{A\gamma\gamma}^{-2}.$ (9)
@@
where $\Phi_{A0}$ is the axion flux at the Earth in case there is no axion decay (6), $\tau_{2\gamma}$ is defined by (8, 9), and $\tau_{f}$ is the time of flight in the axion frame of reference. Because of axion decay, the sensitivity of experiments searching solar axions drops off for large values of $g_{A\gamma}^{2}m_{A}^{3}$. The expected axion spectra for KSVZ- and DFSZ-models as function of $m_{A}$ are shown in Fig. 1 (lines 4 KSVZ-, and lines 5 and 6 for DFSZ-axions). The fluxes have a bell-shaped form with maximum at $\sim(40-50)$ keV since at low $m_{A}$ the probability of decay is low and at large ones the axion flux is reduced due to decays $\rm{A}\rightarrow 2\gamma$ on the way to the Earth.
The results were interpreted in terms of QCD-axion models using relevant relations between the axion mass and the coupling constants. In case of axion-like particles, the QCD-axion relations no longer apply and the results must be analyzed with axion mass and couplings as independent parameters.
@@ 3.2 Compton conversion $A+e\rightarrow\gamma+e$ and axio-electric effect $A+e+Z\rightarrow e+Z$
Two other studied reactions - the Compton process ${\rm A}+e\rightarrow e+\gamma$, and the axioelectric effect ${\rm A}+e+Z\rightarrow e+Z$ - depend on the coupling constant between the axion and the electron $g_{Ae}$.
[…]
@@
$g_{Ae}=\frac{3\alpha^{2}m_{e}}{4\pi^{2}f_{A}}\left(\frac{E}{N}\ln\frac{f_{A}}{m_{e}}-\frac{2}{3}\frac{(4+z)}{(1+z)}\ln\frac{\Lambda}{m_{e}}\right)$ (12)
@@
[…]
@@ 4.1 Borexino detector
The design and main components and features of the Borexino is described in detail in Ago19 ; Ali02 ; Ali09 ; Bac12 ; Bel14 . Borexino is a scintillator detector with an active mass of 278 tons of pseudocumene, doped with $1.5$ g/L of PPO (2.5-diphenyloxazole, a fluorescent dye). The scintillator is housed in a thin nylon inner vessel (IV) and is surrounded by two concentric pseudocumene buffers ($323$ and $567$ tons) doped with a small amount of light quencher to reduce their scintillation. The two buffers are separated by a second thin nylon membrane to prevent diffusion of external radon. The scintillator and buffers are contained in a stainless steel sphere (SSS) with diameter $13.7$ m. The SSS is enclosed in domed water tank (WT), containing 2100 tons of ultra pure water as a shield against external $\gamma$’s and neutrons. The scintillation light is detected by $2212$ $8^{\prime\prime}$ PMTs (PMT - photomultiplier) distributed on the SSS. The WT is equipped with $208$ additional PMTs that act as a Cerenkov muon detector (outer detector) to identify cosmogenic muons. The energy and spatial resolution of the detector were studied with radioactive sources placed at different points inside the IV Bel10 ; Bel10A ; Bel11d .